Circular No. 7021

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

CATACLYSMIC VARIABLES IN PEGASUS M. Modjaz, E. Halderson, T. Shefler, J. Y. King, M. Papenkova, W. D.

Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of a variable star (probably a cataclysmic variable) in our own galaxy during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The variable star was found and confirmed on unfiltered images taken on Sept. 24.3 UT (mag ∼ 18.4) and 25.3 UT (mag ∼ 16.3), respectively. The object is located at α = 23^{h}54^{m}39^{s}.19, δ = +28^{o}22′49′′.7 (equinox 2000.0). A KAIT image of the same field obtained on Sept. 21.3 shows nothing at the position of the variable star (limiting mag ∼ 19.2), nor does the Digitized Sky Survey image of the field shows anything in this position. Spectroscopic classification for the object is encouraged. A finding

chart may be found at http://astron.berkeley.edu/~bait/cv.gif.

X.-y. Dong, J.-y. Wei, Y.-l. Qiu, Q.-y. Qiao, and J.-y. Hu, Beijing Astronomical Observatory (BAO), report their discovery in the course of the BAO supernova survey of another outburst of the cataclysmic variable noted on IAUC 6982. The unfiltered CCD magnitudes of the star were 15.4 on Sept. 22.5 and 17.4 on Sept. 24.5 UT. The star was not present on a CCD image obtained on Sept. 20.

XTE J1946+274 G. L. Israel, Osservatorio Astronomico di Roma; V. F. Polcaro, Isti-

tuto di Astrofisica Spaziale, CNR, Rome; and S. Covino, Osservatorio Astronomico di Brera, report: “A 20-min slitless multi-object spectroscopic image (range 550–800 nm, resolution 8 nm) around the position of the transient x-ray pulsar XTE J1946+274 (IAUC 7014, 7016) was obtained with the 1.5-m telescope (+ BFOSC) at the Astronomical Observatory of Loiano on Sept. 16.92 UT. An Hα emission line (FWHM = 1.2 nm; EW = 0.13 nm) was clearly detected from a star located at α = 19^{h}45^{m}39^{s}.4, δ = +27^{o}22′43′′ (equinox 2000.0). Additional weaker emission lines, possibly from He I, are present. From CCD photometry calibrated by Guide Star Catalog magnitudes, we obtain V = 14.3 ± 0.3 on Sept. 16.8. These findings strongly suggest that this emitting star could be the optical counterpart of XTE J1946+274. Further optical observations of the candidate star are encouraged.”

1998 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green